OpenStudy (anonymous):

What is the 24th term of the arithmetic sequence 10, 16, 22, … ?

OpenStudy (nenadmatematika):

=10+23*6=10+138=148

OpenStudy (anonymous):

you can use the formula an=a + (n-1) d

OpenStudy (anonymous):

a (24th) = 10 + 6*23 = 148

OpenStudy (anonymous):

Thank you so much guys. Sofor this one, would I do the same? What is the 19th term of the arithmetic sequence 15, 23, 31, … ?

OpenStudy (nenadmatematika):

OK try this one by yourself....:D

OpenStudy (anonymous):

Okay :) I am going to try it ... one sec.

OpenStudy (anonymous):

is 151 right?

OpenStudy (anonymous):

a(19th) = 15 + 8*18 = 159

OpenStudy (nenadmatematika):

no moon55 lets do it together...:D First, tell me what is the first element of this sequence?

OpenStudy (anonymous):

10?

OpenStudy (nenadmatematika):

15,23,31,....what is the first number you see?

OpenStudy (anonymous):

oh sorry 15!

OpenStudy (nenadmatematika):

OK that is the first element of your sequence...it's usually marked as a1...OK? now tell me what are the differences between any two consecutive numbers? 15 and 23, 23 and 31 31 and 39?

OpenStudy (anonymous):

By the way, thank you so much for taking the time to explain it like this... and 8?

OpenStudy (nenadmatematika):

you're welcome :D OK it's 8 and the difference is usually marked as ''d'' so in this case your d=8 OK? now when you know the first element of the sequence and difference you have everything to calculate any term....Write down this very important formula for findind any element of the sequence: